# Towards a General-Purpose Zero-Shot Synthetic Low-Light Image and Video Pipeline

### v1

@@ 3.2 Degradation Estimation Network (DEN)
The architecture of the DEN is built upon a U-Net framework [26], with an additional multi-layer perceptron (MLP) [34] head following the encoder. The MLP head is tasked with estimating the noise vector for simulating the noise in the reference data, whereas the decoder reconstructs the denoised version of the input video. The reasoning behind this design is to enable the encoder to effectively discriminate noise from the underlying image content, guiding the MLP head to precise noise vector predictions.
@@
$\mathcal{L}_{rec}=\frac{1}{n}\sum_{i=1}^{n}\lvert\mathbf{x}_{d}-\mathbf{\hat{x}}_{d}\rvert,$ (8)
@@
$\mathcal{L}_{DEN}=\lambda_{1}\mathcal{L}_{MLP}+\lambda_{2}\mathcal{L}_{rec}$ (9)
@@
3.3 Noise Simulator $f(\cdot)$
@@
Most of the noise types used in this simulation are signal-independent and modeled as additive noise, while shot noise is signal-dependent and is typically applied onto the image. However, by adopting the heteroscedatic noise model [19], read-shot noise can be approximated as additive, allowing for easier calculations. As a result, the final simulated noise map is defined as the summation of all individual noise components, expressed as
$f(\mathbf{v})=\mathbf{n}_{h}+\mathbf{n}_{q}+\mathbf{n}_{b}+\mathbf{n}_{b_{t}}+\mathbf{n}_{p},$ (10)
@@
4 Experimental Results and Discussion
4.1 Implementation Details
We utilize the YouTube-VOS dataset [3] for training our network, which includes 3,471 videos representative of diverse real-world scenarios. It also features extensive variability in both object types and movement dynamics for greater generalizability. We train our method for 10 epochs with a batch size of 2 and 16 frames per video. We employ the Adam optimizer with a learning rate of 0.0002, $\beta_{1}$ of 0.5 and $\beta_{2}$ of 0.999. Weights $\lambda_{1}$ and $\lambda_{2}$ are both set to 1.
For performance benchmarking, we compare our framework against several noise models, including Additive White Gaussian Noise (AWGN), Poisson-Gaussian noise (P-G) as approximated by Foi et al. [19], and the Starlight noise generator proposed by Monakhova et al. [22]. The Starlight model is fully re-trained under identical settings as ours and adapted for processing sRGB videos instead of RAW inputs. The consistent training strategy ensures a fair comparison. Importantly, we do not compare our model against other sRGB deep-learning pipelines, such as [27, 28], as they require camera metadata, which would otherwise result in an unfair comparison, and this information is unknown in many real-world cases.
TABLE I: Quantitative Results of Synthetic Pipelines Across Multiple Experiments. The best performers are highlighted in bold, and the second best are underlined.
@@ 4.1 Implementation Details
AWGN 0.458 258.635 0.450 29.716 0.849 0.165 0.205 6.234 18.474 20.328 19.812
P-G [19] 0.358 254.760 0.444 30.222 0.855 0.134 0.199 5.943 25.213 29.167 27.083
Starlight [22] 0.444 252.610 0.428 30.006 0.860 0.143 0.260 5.812 25.337 28.465 28.465
Ours 0.347 211.219 0.337 30.272 0.864 0.130 0.200 5.565 29.919 32.343 31.291
4.2 Synthetic Quality Assessment
For a quantitative analysis of the performance of each synthetic pipeline, we calculate the Kullback–Leibler divergence (KLD), Fréchet Inception Distance (FID) [35] and Kernel Inception Distance (KID) [36] between the real noise maps and the synthetic noise maps. KL divergence is a broadly-applied statistical measure for quantifying the distance between probability distributions, and therefore is a highly suitable metric for our assessment. FID and KID are two popular metrics used in image generation tasks for measuring the performance of generative models. FID measures the distance between two multivariate Gaussians fitted to feature representations of the input dataset against those of the target set. Unlike FID, KID does not assume a Gaussian distribution but instead uses Maximum Mean Discrepancy.
We use the SIDD-Small dataset [37] for evaluation as it covers a wide range of noise characteristics, which are influenced by various ISO levels, shutter speeds, illuminant temperatures, lighting conditions) across 5 different camera models. Moreover, the clean images are provided, allowing for the computation of real noise maps. The results, presented in Table I (under ‘Synthetic Quality’), demonstrate that our synthetic noise pipeline achieves the best results for all 3 metrics relative to the real noise, indicating a higher fidelity in noise replication. We would like to note that the FID scores are inflated, reflecting the metric’s strong bias with respect to sample size.
We also present visual comparisons of the results from each synthetic pipeline in Fig. 4 for qualitative analysis. Our method exhibits superior performance, adapting noise parameters dynamically to the noisy reference input. Subjectively, our results appear closer to the reference (noted as ‘Real’ in the figure) and do not overestimate noise like other methods. However, as shown in Fig. 4, there is evident noise reduction applied in the real noisy images, but our pipeline does not incorporate blur effects or other in-camera post-processing. These complex degradations will be addressed in our future work.
@@ 4.2 Synthetic Quality Assessment
Figure 4: Visual comparison of different synthetic pipelines compared against the real low-light images from the SIDD-Small dataset [37]. The KL divergence for each synthetic image is also provided.
4.3 Low-Light Video Enhancement (LLVE)
@@
Figure 5: Qualitative comparison of the outputs of four BVI-Mamba [38] models trained with their respective noise pipelines on an extreme example of low-light noise. Simpler models are unable to capture the high degree of variability in the scene’s noise, while more sophisticated models such as Starlight and ours allow BVI-Mamba to learn to denoise more effectively.
@@
We assess the performance of our noise simulator on low-light video enhancement through quantitative and qualitative analysis. Given that most low-light enhancement methods are proposed for images, and because there are fewer video-specific methods due to the increased computational costs, we select a pre-trained low-light video enhancement model BVI-Mamba [38], which reduces computational complexity by employing Visual State Space (VSS) blocks. The dataset used for pre-training is the BVI-RLV [38] dataset, which offers over 30K fully registered HD paired video frames in normal and dark conditions with predefined training and testing video sets.
We conduct testing using a real noisy low-light HD video, referred to as the ‘horse’ video, captured outdoors after sunset with a Canon ML-105 camera (Fig. 5 left). Our synthetic noise is generated using the DEN model, with this ‘horse’ video serving as $\mathbf{y}_{n}$. For training, synthetic noise from each pipeline is added to the BVI-RLV training set following Equation 1.
@@
The first experiment analyzes how well the low-light enhancer can learn different noise characteristics. We independently apply the same synthetic noise from each pipeline to the BVI-RLV testing sets. The enhanced outputs are then compared with the clean, normal-light ground truths. Objective metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and perceptual similarity metrics like LPIPS [39], are used to evaluate the performance of four different pipelines. The quantitative results can be seen in Table I.
@@
The second analysis evaluates how closely the noise simulators can replicate the characteristics of the ’horse’ video. We enhance this real video using models trained with various synthetic noises. As there is no ground truth available, we calculate no-reference metrics CLIP Image Quality Assessment (CLIP-IQA) [40] and Naturalness Image Quality Evaluator (NIQE) [41], for a quantitative comparison. Higher CLIP-IQA scores and lower NIQE scores indicate better perceptual quality. The prompts used for CLIP-IQA are ‘quality’, ‘brightness’, and ‘noisiness’. Visual comparisons are shown in Fig. 5. Models trained with synthetic noise from our pipeline achieve a better balance between perceptual quality and naturalness. In smooth, bright regions such as the sky, shown in the top row, our method achieves the best NIQE scores compared to all other pipelines. Although Starlight has the highest CLIP-IQA, it produces artifacts like color distortion. In textured, dark areas, such as the tree region shown in the bottom row, Starlight and P-G achieve better CLIP-IQA scores, suggesting they might be better at enhancing brightness or details in those regions. However, results trained on our pipeline still maintain competitive NIQE scores.
We further evaluate the impact of enhancement on object detection performance by using the pre-trained YOLOv11n model to detect a horse located in the middle of a scene. Importantly, this object was not detectable in the original noisy video due to the provided model being trained only on clean images. Post-enhancement results, depicted in Fig. 6, reveal that the YOLOv11n model successfully detects the horse in the output enhanced by BVI-Mamba trained on our synthetic data. In contrast, the horse is mislabeled when using the enhancer trained on Starlight data. When enhanced with AWGN data, the model erroneously predicts the presence of both a cow and a horse. Although the P-G data allows for successful horse detection, the confidence of the YOLOv11n model is notably lower compared to its performance with our synthetic data.
@@
Figure 6: Qualitative comparison of the pre-trained YOLOv11n [42] on outputs of four BVI-Mamba models trained with their respective noise pipelines. The denoiser trained using our approach allows successful detection of a horse, while the model trained using Starlight results in the detection of a cow.
4.4 Object Detection
(a) Pre-trained YOLOv11n on noisy input.
(b) Our YOLOv11n on noisy input.
(c) Pre-trained YOLOv11n on clean input.
Figure 7: Qualitative comparison showing the benefits of using our approach to train an object detection model for use in low-light conditions. Our model closely matches the performance of the pre-trained model on the clean scene, without the need for a denoising pre-processing step.
@@ 4.4 Object Detection
To determine the effectiveness of our synthetic pipeline for downstream tasks separate to video enhancement, we conduct object detection experiments under real low-light conditions. We select the smallest model, ‘nano’, of the ubiquitous YOLOv11 [42] as our test model, due to its excellent performance-to-speed ratio. For each of the four different synthetic pipelines, we train a YOLOv11n model on the popular Microsoft COCO dataset [43]. The synthetic noise is inserted after simulating low-light conditions, using Equation 1. Each model is trained on its respective synthetic low-light data for 100 epochs, and the best hyperparameters are selected from the validation split, which has the same synthetic noise type as the training dataset.
@@
The models are evaluated on the highest ISO Sony frames from the real low-light image dataset [44], which exhibits higher noise levels relative to the other existing real low-light datasets. The performance of each model is tested using pseudo-ground-truth annotations generated by applying the pre-trained YOLOv11n onto the clean ISO-100 videos. In 4 of the 20 scenes, the pre-trained model does not detect any instances of the 80 COCO classes, so those scenes are removed from the test set. We measure this performance using the standard object detection metric of Average Precision (AP), over various Intersection over Union (IoU) thresholds. Specifically, we calculate AP50-95, in steps of 5, AP50, and AP75.
@@
As shown in Table I (under ‘Object Detection’), training the YOLOv11n model using our training approach results in improved precision across all the thresholds. This demonstrates that our synthetic noise most closely resembles the characteristics of the real low-light test set and enhances the noise handling and feature extraction capabilities of the object detector more effectively. Qualitative results are shown in Fig. 7, along with an example of the limitations of applying a clean model to noisy input. Without training directly on noisy images, the pre-trained YOLOv11n struggles to differentiate cups and bowls, and ignores the presence of a dining table. However, training using our synthetic noise, sampled from the noise vectors predicted by the DEN on the dataset, results in much greater performance.
@@
It is noted that the pre-trained YOLOv11n applied to the clean image is not without fault. It fails to detect the other four bowls in the image with high confidence, and only exhibits particularly high confidence for the cups in the scene. In addition, these three cases exhibit higher performance due to the relatively high light conditions.
4.5 Ablation Study
We characterize the significance of each component in our network through an ablation study. We conduct 3 experiments, each with a different combination of the components, calculating KLD, FID and KID of the noise maps for each experiment. For model v1 in the ablation study, we simply train a U-Net [26] to generate a synthetic noisy image, without employing $f(\cdot)$. This model predictably performs the worst, as the network cannot easily learn to generate appropriate noise characteristics using only a clean input and training data of varying noise characteristics without additional physics-based guidance. For model v2, we consider a network architecture with an encoder and an MLP head. This network was trained to directly estimate the noise vector used to synthesize the noise characteristics from the reference input. This iteration showed some improvements by mostly generating plausible noise, but failed to capture the exact features of the desired noise. Model v3 is the network architecture that we propose for the DEN, comprised of an encoder, with both MLP and decoder heads. This method improves upon model v2 due to the added decoder head and accompanying $\mathcal{L}_{rec}$, forcing the model to learn the specifics of the reference noise.

### v2

@@ 3.2. Degradation Estimation Network (DEN)
The architecture of the DEN is built upon a U-Net framework (Ronneberger et al., 2015), with an additional multi-layer perceptron (MLP) (Rosenblatt, 1958) head following the encoder. The MLP head is tasked with estimating the noise vector for simulating the noise in the reference data, whereas the decoder reconstructs the denoised version of the input video. The reasoning behind this design is to enable the encoder to effectively discriminate noise from the underlying image content, guiding the MLP head to precise noise vector predictions.
@@
(8) $\mathcal{L}_{rec}=\frac{1}{n}\sum_{i=1}^{n}\lvert\mathbf{x}_{d}-\mathbf{\hat{x}}_{d}\rvert$,
@@
(9) $\mathcal{L}_{DEN}=\lambda_{1}\mathcal{L}_{MLP}+\lambda_{2}\mathcal{L}_{rec},$
@@
3.3. Noise Simulator $f(\cdot)$
@@
The final simulated noise map is defined as the summation of all individual noise components, expressed as
(10) $f(\mathbf{v})=\mathbf{n}_{h}+\mathbf{n}_{q}+\mathbf{n}_{b}+\mathbf{n}_{b_{t}}+\mathbf{n}_{p},$
@@
4. Experimental Results and Discussion
4.1. Implementation Details
We utilize the YouTube-VOS dataset (Xu et al., 2018) for training our network, which includes 3,471 videos representative of diverse real-world scenarios. It also features extensive variability in both object types and movement dynamics for greater generalizability. We train our method for 10 epochs with a batch size of 2 and 16 frames per video. We employ the Adam optimizer with a learning rate of 0.0002, $\beta_{1}$ of 0.5 and $\beta_{2}$ of 0.999. Weights $\lambda_{1}$ and $\lambda_{2}$ are both set to 1.
For performance benchmarking, we compare our framework against several noise models, including Additive White Gaussian Noise (AWGN), Poisson-Gaussian noise (P-G) as approximated by Foi et al. (Foi et al., 2008), and the Starlight noise generator proposed by Monakhova et al. (Monakhova et al., 2022). The Starlight model is fully re-trained under identical settings as ours and adapted for processing sRGB videos instead of RAW inputs. The consistent training strategy ensures a fair comparison. Importantly, we do not compare our model against other sRGB deep-learning pipelines, such as (Kousha et al., 2022; Fu et al., 2023a), as they require camera metadata, which would otherwise result in an unfair comparison, and this information is unknown in many real-world cases.
Table 1. Quantitative Results of Synthetic Pipelines Across Multiple Experiments. The best performers are highlighted in bold, and the second best are underlined.
@@ 4.1. Implementation Details
AWGN 0.458 258.635 0.450 29.716 0.849 0.165 0.777 6.234 18.474 20.328 19.812
P-G (Foi et al., 2008) 0.358 254.760 0.444 30.222 0.855 0.134 0.763 5.943 25.213 29.167 27.083
Starlight (Monakhova et al., 2022) 0.444 252.610 0.428 30.006 0.860 0.143 0.827 5.812 25.337 28.465 28.465
Ours 0.347 211.219 0.337 30.272 0.864 0.130 0.775 5.565 29.919 32.343 31.291
4.2. Synthetic Quality Assessment
For a quantitative analysis of the performance of each synthetic pipeline, we calculate the Kullback–Leibler divergence (KLD), Fréchet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) between the real noise maps and the synthetic noise maps. KL divergence is a broadly-applied statistical measure for quantifying the distance between probability distributions, and therefore is a highly suitable metric for our assessment. FID and KID are two popular metrics used in image generation tasks for measuring the performance of generative models. FID measures the distance between two multivariate Gaussians fitted to feature representations of the input dataset against those of the target set. Unlike FID, KID does not assume a Gaussian distribution but instead uses Maximum Mean Discrepancy and is more suitable for small dataset sizes.
We use the SIDD-Small dataset (Abdelhamed et al., 2018) for evaluation as it covers a wide range of noise characteristics (which are influenced by various camera settings) across 5 different camera models. Moreover, the clean images are provided, allowing for the computation of real noise maps. For evaluation, we selected only the low-light scenes. The results, presented in Table 1 (under ‘Synthetic Quality’), demonstrate that our synthetic noise pipeline achieves the best results for all 3 metrics relative to the real noise, indicating a higher fidelity in noise replication. We would like to note that the FID scores are inflated, reflecting the metric’s strong bias to sample size.
We also present visual comparisons of the results from each synthetic pipeline in Fig. 3 for qualitative analysis. Our method exhibits superior performance, adapting noise parameters dynamically to the noisy reference input. Subjectively, our results appear closer to the reference (noted as ‘Real’ in the figure) and do not overestimate noise like other methods. However, as shown in Fig. 3, there is evident noise reduction applied in the real noisy images which is not considered in our pipeline. These complex degradations will be addressed in our future work.
@@ 4.2. Synthetic Quality Assessment
Figure 3. Visual comparison of different synthetic pipelines compared against the real low-light images from the SIDD-Small dataset (Abdelhamed et al., 2018). The KL divergence for each synthetic image is also provided.
[…]
Examples of applying various synthetic noise pipelines onto patches from the SIDD-Small dataset, along with the KL divergence values, with ours visually appearing the most similar to the ground truth for every example.
4.3. Low-Light Video Enhancement (LLVE)
@@
Figure 4. Qualitative comparison of the outputs of four BVI-Mamba (Lin et al., 2024a) models trained with their respective noise pipelines on an extreme example of low-light noise. Simpler models are unable to capture the high degree of variability in the scene’s noise, while more sophisticated models such as Starlight and ours allow BVI-Mamba to learn to denoise more effectively.
[…]
Example patches on a frame from the horse video, showing the denoising performances using BVI-Mamba trained on the various synthetic pipelines.
@@
We assess the performance of our synthetic noise pipeline for low-light video enhancement through quantitative and qualitative analysis. For the low-light video enhancement experiments, we select a pre-trained model, BVI-Mamba (Lin et al., 2024a), as there are a limited number of video-specific LLVE methods. BVI-Mamba employs Visual State Space (VSS) blocks to reduce computational complexity. BVI-Mamba was pre-trained on the BVI-RLV (Lin et al., 2024a) dataset, with over 30K fully-registered HD paired frames in normal/low-light conditions. We conduct our LLVE evaluation on a real noisy low-light video, which we refer to as the ‘horse’ video. The ‘horse’ video is captured outdoors after sunset with a Canon ML-105 camera (Fig. 4 left). For our synthetic pipeline, we first passed the horse video into DEN to estimate the noise parameters $\mathbf{v}_{horse}$, which is then used to generate the noisy videos for finetuning.
@@
We first analyze how well the pre-trained low-light enhancer can learn different noise characteristics. We independently apply the same synthetic noise from each pipeline to the BVI-RLV testing sets and pass these synthesized inputs into the pre-trained BVI-Mamba model. The enhanced outputs are then compared with the clean, normal-light ground truths. Objective metrics include Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and perceptual similarity metrics like Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), are used to evaluate the performance of four different pipelines. The quantitative results can be seen in Table 1. The second analysis evaluates how closely the noise simulators can replicate the characteristics of the ‘horse’ video. We train various BVI-Mamba models on different synthetic noise pipelines and evaluate their enhancing performances on the ‘horse’ video. For our method, we first pass the ‘horse’ video into DEN to estimate the noise parameters $\mathbf{v}_{horse}$, which are then used for synthesizing the noisy training data. As there is no clean ground-truth available for the ‘horse’ video, we calculate no-reference metrics CLIP Image Quality Assessment (CLIP-IQA) (Wang et al., 2023) and Naturalness Image Quality Evaluator (NIQE) (Mittal et al., 2013), for a quantitative comparison. Higher CLIP-IQA scores and lower NIQE scores indicate better perceptual quality. The prompts used for CLIP-IQA are ‘quality’, ‘brightness’, and ‘noisiness’. Visual comparisons are shown in Fig. 4.
@@
Models trained with synthetic noise from our pipeline achieve a better balance between perceptual quality and naturalness. In smooth, bright regions, such as the sky shown in the top row, our method achieves the best NIQE scores compared to all other pipelines, while Starlight produces artifacts like color distortion. Although P-G has the highest CLIP-IQA score, our result visually exhibits less noise. In textured, dark areas, such as the tree region shown in the bottom row, P-G achieves better CLIP-IQA and NIQE scores, suggesting it might be better at enhancing brightness in those regions. However, results trained on our pipeline still maintain a competitive balance between detail restoration and denoising. We also compare enhancement performances on starlight videos from (Monakhova et al., 2022), with BVI-Mamba trained on our synthetic pipeline shows competitive results 5.
We further evaluate the enhancing capabilities of each synthetic pipeline by showing the improvements to object detection performances after enhancing the ‘horse’ video. For this experiment, we use a pre-trained YOLOv11n model for detecting the horse in the video. Without any enhancement, the detection model was unable to detect the horse, which is expected as it was not trained on noisy data. After passing the ‘horse’ video into the BVI-Mamba model trained with our synthetic pipeline, the pre-trained YOLOv11n model successfully detected the horse in the scene. In comparison, the BVI-Mamba model trained on the Starlight pipeline performed the worst by mislabeling the horse as a cow, while the model trained on AWGN noise resulted in two detections; a horse and a cow. Although the P-G data allows for successful horse detection, the confidence score is notably lower compared to its performance with our synthetic data.
Raw
Starlight (Monakhova et al., 2022)
Ours
Figure 5. Qualitative comparison of the outputs of BVI-Mamba models trained with receptive noise pipelines on an example patch from the Submillilux dataset (Monakhova et al., 2022).
[…]
The denoising performances of the BVI-Mamba model trained on the different noise pipelinesn and tested on the Submillilux dataset.
@@
Figure 6. Qualitative comparison of the pre-trained YOLOv11n (Jocher et al., 2023) on outputs of four BVI-Mamba models trained with their respective noise pipelines. The denoiser trained using our approach has the most successful result.
[…]
The detection performances of a pre-trained YOLOv11n model on the denoised outputs from the BVI-Mamba models trained on the various noise pipelines.
4.4. Object Detection
Figure 7. Qualitative comparison showing the performance improvements when using our synthetic noise pipeline to train an object detection model for use in low-light conditions on an image from BVI-Lowlight dataset (Anantrasirichai et al., 2024). Left: pre-trained YOLOv11n (Jocher et al., 2023). Right: YOLOv11n trained with our synthetic data.
[…]
Comparison photos of object detection performances using a pre-trained YOLOv11n model and a YOLOv11n model trained on our synthetic data.
@@ 4.4. Object Detection
To determine the effectiveness of our synthetic pipeline for downstream tasks separate to video enhancement, we conduct object detection experiments under real low-light conditions. For each of the four different synthetic pipelines, we train a YOLOv11n (Jocher et al., 2023) model on the popular Microsoft COCO dataset (Lin et al., 2014). The synthetic noise is inserted after simulating low-light conditions, using Equation 1. Each model is trained on its respective synthetic low-light data for 100 epochs.
@@
The models are evaluated on the highest ISO frames from the real low-light image dataset (Anantrasirichai et al., 2024), which exhibits higher noise levels relative to the other existing real low-light datasets. The performance of each model is tested using pseudo-ground-truth annotations generated by applying the pre-trained YOLOv11n onto the clean ISO-100 videos. In 4 of the 20 scenes, the pre-trained model did not detect any instances of the 80 COCO classes, so those scenes are removed from the test set. We measured this performance using the standard object detection metric Average Precision (AP), over various Intersection over Union (IoU) thresholds; AP50-95, AP50, and AP75.
@@
As shown in Table 1 (under ‘Object Detection’), training the YOLOv11n model using our training approach results in improved precision across all thresholds. This demonstrates that our synthetic noise is capable of reproducing the noise characteristics of the dataset, and suitable for low-light domain adaptation. Examples of detection results are shown in Fig. 7; without training directly on noisy images, the pre-trained YOLOv11n struggled to differentiate cups and bowls, and ignores the presence of a dining table. However, training on our synthetic pipeline, resulted in much greater performance, with correct classification of the cups and less false positive detections.
@@
It is noted that the pre-trained YOLOv11n applied to the clean image (for pseudo ground-truth) is not without fault. It failed to detect the other four bowls in the image with high confidence, and only exhibited particularly high confidence for the cups in the scene. In addition, these three cases exhibited higher performance due to the relatively high light conditions.
4.5. Ablation Study
We assess the significance of each component in our network architecture through an ablation study, showing the results in Table 2. We conduct 3 experiments, each with a different combination of the components, calculating KLD, FID and KID of the noise maps for each experiment. We refer to the different architectures as model v1, v2, and v3. Model v1 is simply a U-Net (Ronneberger et al., 2015) trained to generate a synthetic noisy image without employing $f(\cdot)$. Model v2 resembles a classification network, with an encoder and a MLP head for estimating the noise parameters. Model v3, the proposed architecture for DEN, is comprised of an encoder, a decoder and an MLP head, allowing for the network to understand the noise characteristics further via the dual network heads.